jim has gotten scores of 99 and 93 on his firs tests. what score must he grt on his third test to keep an average of 90 or greater
Answer by friesr(113)

Jim will need a total of 90 * 3 = 270 point.
-
Currently Jim has 192 points.
-
270 - 192 = 78
-
Jim will need to make a 78 on his third test
===
99 + 93 + 78 = 270 / 3 = 90